# Projective rigidity of point-line configurations in the plane

### v1

@@ 6. Symmetric projective configurations
6.1. Symmetries, dualities and polarities of projective space
@@
Figure 10. An autopolar configuration: the polarity $\pi_{C}$ sends the configuration to itself, i.e. it sends the points of the configuration to the lines of the configuration and vice versa.
@@
6.2. Projective orbit rigidity matrix
@@
Throughout this section, we will let $\Gamma$ denote a subgroup of the group of correlations of a projective configuration such that each element of $\Gamma$ is either an element of $PO(3)$ or a polarity defined by an orthogonal matrix. We say that an infinitesimal projective motion of the configuration is “$\Gamma$-symmetric” if it also has symmetry $\Gamma$. More formally, let $m$ be an element of the kernel of the projective rigidity matrix and let $\gamma$ be an element of $\Gamma$. The vector $m$ has two entries for each point of the configuration, and two entries for each line of the configuration. Let $m(p)$ denote the vector consisting of the two elements corresponding to the point $p$. We can then require that $m(\gamma p)=\gamma m(p)$ for all points $p$ of the configuration.
@@ 6.2.1. Free actions
(9) $q\cdot\gamma m(r)+\gamma r\cdot m(q)=0.$
@@
(10) $\gamma^{-1}q\cdot m(r)+\gamma r\cdot m(q)=0.$
@@
Theorem 6.1.
Let $\mathbf{p}$ be a configuration of points and lines with symmetry group $\Gamma$. Then $m$ is an element of the kernel of $M_{2}^{\Gamma}(S,\mathbf{p})$ if and only if it is the restriction of a $\Gamma$-symmetric infinitesimal projective motion of the configuration.
@@
(12) $\begin{split}\beta q\cdot\beta\gamma\bar{m}(r)+\beta\gamma r\cdot\beta\bar{m}(q)=0.\end{split}$
@@
6.2.2. Actions with fixed points, lines, and incidences
@@
(13) $q\cdot\gamma M_{r}\hat{m}(r)+\gamma r\cdot M_{q}\hat{m}(q)=0$
@@
(14) $\gamma^{-1}q\cdot M_{r}\hat{m}(r)+\gamma r\cdot M_{q}\hat{m}(q)=0$
@@
Theorem 6.2.
Let $\mathbf{p}$ be a $\Gamma$-symmetric configuration of points and lines. Then $\hat{m}$ is an element of the kernel of $M^{\Gamma}(S,\mathbf{p},\mathbf{l})$ if and only if $\bar{m}$ defined by $\bar{m}(r)=M_{r}\hat{m}(r)$ for each representative $r$ of the point and line orbits under the action of $\Gamma$ is the restriction of a $\Gamma$-symmetric infinitesimal projective motion of $S$.
@@
Suppose that $(q,\gamma r)\in I$ is a representative of an orbit of incidences.
The equation in $M^{\Gamma}(S,\mathbf{p},\mathbf{l})$ corresponding to $(q,\gamma r)$ is Equation (14).
@@
(16) $\begin{split}\beta q\cdot\beta\gamma M_{r}\hat{m}(r)+\beta\gamma r\cdot\beta M_{q}\hat{m}(q)=0.\end{split}$
@@
Figure 11. A configuration with dihedral symmetry $D_{4}$.
@@
Figure 12. A non-trivial motion of the configuration in Figure 10 preserving autopolarity.
@@
Under the polarity $\pi_{C}$, there are six orbits $\{p_{i},L_{i}\}$ of points and lines. There are eight orbits of incidences, namely $i_{0}=\{(p_{1},L_{1})\}$, $i_{1}=\{(p_{1},L_{2}),(p_{2},L_{1})\}$, $i_{2}=\{(p_{1},L_{3}),(p_{3},L_{1})\}$, $i_{3}=\{(p_{4},L_{4})\}$, $i_{4}=\{(p_{4},L_{5}),(p_{5},L_{4})\}$, $i_{5}=\{(p_{4},L_{6}),(p_{6},L_{4})\}$, $i_{6}=\{(p_{2},L_{6}),(p_{6},L_{2})\}$ and $i_{7}=\{(p_{3},L_{5}),(p_{5},L_{3})\}$. The orbit rigidity matrix $M_{2}^{\langle\pi_{C}\rangle}(S,\mathbf{p})$ is
@@
7. Conclusion and future work
@@
Inductive constructions. There are well-known inductive constructions, known as 00- and $1$-extensions, that preserve rigidity of bar-joint frameworks in $\mathbb{R}^{d}$. In the plane, 00- and $1$-extensions are sufficient for inductively constructing all minimally rigid graphs, starting from a single edge. We have shown that operations in the Cayley algebra, i.e. meets and joins, preserve independence in the projective rigidity matrix, and can therefore be used to inductively construct projectively rigid configurations. However, operations in the Cayley algebra are clearly not sufficient for constructing all projectively rigid configurations, so it would be interesting to find more inductive constructions that preserve projective rigidity.
@@
Families of rigid $v_{k}$-configurations. In Example 4.3, we show a projectively rigid $20_{4}$-configuration. As $v_{k}$-configurations are overconstrained with respect to the count in (4) whenever $k\geq 4$, it seems likely that there are more projectively rigid $v_{k}$-configurations. Constructing families of projectively rigid $v_{k}$-configurations, or proving that known families of $v_{k}$-configurations are projectively rigid, is a potential direction for future research.

### v3

@@ 5 Symmetric projective configurations
5.1 Symmetries, dualities and polarities of projective space
@@
Figure 5: An autopolar configuration: the polarity $\pi_{C}$ sends the configuration to itself, i.e. it sends the points of the configuration to the lines of the configuration and vice versa.
@@
5.2 Projective orbit rigidity matrix
@@
Throughout this section, we will let $\Gamma$ denote a subgroup of the group of correlations of a projective configuration such that each element of $\Gamma$ is either an element of $PO(3)$ or a polarity defined by an orthogonal matrix. We say that an infinitesimal projective motion of the configuration is “$\Gamma$-symmetric” if it also has symmetry $\Gamma$.
More formally, let $m$ be an element of the kernel of the projective rigidity matrix and let $\gamma$ be an element of $\Gamma$. The vector $m$ has two entries for each point of the configuration, and two entries for each line of the configuration. Let $m(p)$ denote the vector consisting of the two elements corresponding to the point $p$. We can then require that $m(\gamma p)=\gamma m(p)$ for all points $p$ of the configuration.
@@ 5.2.1 Free actions
$q\cdot\gamma m(r)+\gamma r\cdot m(q)=0.$ (9)
@@
$\gamma^{-1}q\cdot m(r)+\gamma r\cdot m(q)=0.$ (10)
@@
Theorem 4.
Let $(S,\mathbf{p},\mathbf{l})$ be a configuration of points and lines with symmetry group $\Gamma$. Then $m$ is an element of the kernel of $M^{\Gamma}(S,\mathbf{p},\mathbf{l})$ if and only if it is the restriction of a $\Gamma$-symmetric infinitesimal projective motion of the configuration.
@@
$\begin{split}\beta q\cdot\beta\gamma\bar{m}(r)+\beta\gamma r\cdot\beta\bar{m}(q)=0.\end{split}$ (12)
@@
5.2.2 Actions with fixed points, lines, and incidences
@@
$q\cdot\gamma M_{r}\hat{m}(r)+\gamma r\cdot M_{q}\hat{m}(q)=0$ (13)
@@
$\gamma^{-1}q\cdot M_{r}\hat{m}(r)+\gamma r\cdot M_{q}\hat{m}(q)=0$ (14)
@@
Theorem 5.
Let $(S,\mathbf{p},\mathbf{l})$ be a $\Gamma$-symmetric configuration of points and lines. Then $\hat{m}$ is an element of the kernel of $M^{\Gamma}(S,\mathbf{p},\mathbf{l})$ if and only if $\bar{m}$ defined by $\bar{m}(r)=M_{r}\hat{m}(r)$ for each representative $r$ of the point and line orbits under the action of $\Gamma$ is the restriction of a $\Gamma$-symmetric infinitesimal projective motion of $S$.
@@
Suppose that $(q,\gamma r)\in I$ is a representative of an orbit of incidences. The equation in $M^{\Gamma}(S,\mathbf{p},\mathbf{l})$ corresponding to $(q,\gamma r)$ is Equation (14).
@@
$\begin{split}\beta q\cdot\beta\gamma M_{r}\hat{m}(r)+\beta\gamma r\cdot\beta M_{q}\hat{m}(q)=0.\end{split}$ (16)
@@
Figure 6: A configuration with dihedral symmetry $D_{4}$.
@@
Figure 7: A non-trivial motion of the configuration in Figure 5 preserving autopolarity.
@@
Under the polarity $\pi_{C}$, there are six orbits $\{p_{i},L_{i}\}$ of points and lines. There are eight orbits of incidences: $i_{0}=\{(p_{1},L_{1})\}$, $i_{1}=\{(p_{1},L_{2}),(p_{2},L_{1})\}$, $i_{2}=\{(p_{1},L_{3}),(p_{3},L_{1})\}$, $i_{3}=\{(p_{4},L_{4})\}$, $i_{4}=\{(p_{4},L_{5}),(p_{5},L_{4})\}$,$i_{5}=\{(p_{4},L_{6}),$
$(p_{6},L_{4})\}$, $i_{6}=\{(p_{2},L_{6}),(p_{6},L_{2})\}$ and $i_{7}=\{(p_{3},L_{5}),(p_{5},L_{3})\}$. The orbit rigidity matrix $M_{2}^{\langle\pi_{C}\rangle}(S,\mathbf{p})$ is
@@
6 Conclusion and future work
@@
Implications of self-stresses for realizability of incidence geometries. What are the implications of dependencies among the incidences of an incidence geometry for whether or not the incidence geometry is realizable as points and straight lines in the projective plane? Does non-realizability of an incidence geometry imply that the incidences are dependent, or are there non-realizable incidence geometries with sets of incidences that are independent in the projective rigidity matroid? Note that points or lines coinciding does not necessarily imply that the incidences are dependent.
Inductive constructions. There are well-known constructions, known as 00- and $1$-extensions, that preserve rigidity of bar-joint frameworks in $\mathbb{R}^{d}$. In the plane, 00- and $1$-extensions are sufficient for inductively constructing all minimally rigid graphs, starting from a single edge.
Operations in the Cayley algebra, i.e. adding a line between two points not already connected by a line in the configuration, or adding the intersection point of two lines whose point of intersection is not already in the configuration, preserve independence in the projective rigidity matrix, and can therefore be used to inductively construct projectively rigid configurations [6]. However, operations in the Cayley algebra are clearly not sufficient for constructing all projectively rigid configurations, so it would be interesting to find more inductive constructions that preserve projective rigidity.
@@
Families of rigid $v_{k}$-configurations. In Example 2, we show a projectively rigid $20_{4}$-configuration. As $v_{k}$-configurations are overconstrained with respect to the count in (5) whenever $k\geq 4$, it seems likely that there are more projectively rigid $v_{k}$-configurations. Constructing families of projectively rigid $v_{k}$-configurations, or proving that known families of $v_{k}$-configurations are projectively rigid, is another potential direction for future research.
Acknowledgements
Most of our results presented here were obtained during the Focus Program on Geometric Constraint Systems July 1 - August 31, 2023 and we gratefully acknowledge the productive work environment provided by the Fields Institute for Research in Mathematical Sciences, as well as its financial support. The contents of this article are solely the responsibility of the authors and do not necessarily represent the official views of the Institute.
The work was also supported by the Knut and Alice Wallenberg Foundation Grant 2020.0001 and 2020.0007.